# Counterfactuals As a Means for Evaluating Faithfulness of Attribution Methods in Autoregressive Language Models

### v1

@@ Abstract
Despite the widespread adoption of autoregressive language models, explainability evaluation research has predominantly focused on span infilling and masked language models (MLMs). Evaluating the faithfulness of an explanation method—how accurately the method explains the inner workings and decision-making of the model—is very challenging because it is very hard to separate the model from its explanation. Most faithfulness evaluation techniques corrupt or remove some input tokens considered important according to a particular attribution (feature importance) method and observe the change in the model’s output. This approach creates out-of-distribution inputs for causal language models (CLMs) due to their training objective of next token prediction. In this study, we propose a technique that leverages counterfactual generation to evaluate the faithfulness of attribution methods for autoregressive language modeling scenarios. Our technique creates fluent and in-distribution counterfactuals that makes evaluation protocol more reliable.
@@ 1 Introduction
Most state-of-the-art NLP systems use autoregressive transformer-based language models (Touvron et al., 2023; Brown et al., 2020; Groeneveld et al., 2024). Since these models are opaque, there is great motivation to understand their decision-making process, and so, explanation methods are becoming increasingly important.
Attribution or feature importance (FI) is a common approach for explaining models (Zhao et al., 2023). Attribution methods’ objective is to explain which input features are most salient in the model’s predictions. These methods achieve this objective by assigning a scalar to each input feature, representing its significance to the model’s decision-making process. Input features in most NLP tasks are sub-words or some combination of them.
@@
A pressing issue for testing and evaluating faithfulness of attribution methods is that most techniques focus on denoising language models or masked language models (Kobayashi et al., 2020, 2021; Ferrando et al., 2022b; Modarressi et al., 2022, 2023; Mohebbi et al., 2023), and most recent work dealing with autoregressive language models has focused on the plausibility of attributions (Yin and Neubig, 2022; Ferrando et al., 2023). Plausibility (or persuasiveness) might be an objective of the explainer, but the explainee’s objective is to understand the model’s decision-making process, not being persuaded that the decision is correct (Jacovi and Goldberg, 2021).
Almost all previous faithfulness evaluation techniques corrupt the input in one way or another, i.e., masking or erasing important tokens according to the attribution technique, and then looking at the change in the prediction. These techniques probably work fine for MLMs, which are pre-trained for mask and span infilling. However, for autoregressive language models like GPT-2, pre-trained for next token prediction, masking or erasing creates an out-of-distribution (OOD) input for the model. In this case, it is unclear whether corruption techniques evaluate the informativeness of the corrupted tokens or the robustness of the model to unnatural text and the artifacts introduced during test time (Hooker et al., 2019). Furthermore, inputs being OOD makes the explanation socially misaligned Hase et al. (2021); meaning what people expect from FI explanations—demonstrating which features are the most important in the model’s decision-making—and what the FI shows become different because FI becomes dependent on model priors rather than removal of features that the model has learned to rely on.
In this work, inspired by counterfactual generation—changing the model’s input in a way that flips the output—we develop a technique for evaluating the faithfulness of attribution methods in the autoregressive generation scenario. We use counterfactual generators to change the input, focusing on the important tokens specified by the attribution methods, and make sure that the input to the model is natural, fluent, and in-distribution. This way, we know that the change in the model’s prediction is because of changing the important tokens and not because of the input being out of distribution.
We argue that if an attribution method helps the counterfactual generator to change the model’s prediction with fewer changes, that method knows more about the model’s inner workings, which means it is more faithful. Also, because of the large output space of autoregressive language models like GPT-2 and LLaMA, including often thousands of vocabulary items, looking at the entire output space does not provide much insight. We use contrastive explanation Yin and Neubig (2022), which means looking only at the token predicted by the model and a foil token.
@@
We apply our faithfulness evaluation approach to several attribution techniques including gradient norm, gradient $\times$ input, Erasure, KernelSHAP, and integrated gradient in the next word prediction of two LMs: our fine-tuned Gemma-2b and off-the-shelf Gemma-2b-instruct (Team et al., 2024).
@@
Our contributions are as follows: (i) We introduce a novel faithfulness evaluation protocol, focused on maintaining the input data distribution of the model, suitable for the attribution methods of language models. (ii) We evaluate and rank several popular attribution methods using our approach. (iii) We demonstrate the difference in sensitivity between fine-tuned and off-the-shelf models when dealing with out-of-distribution (OOD) data and provide a solution.
@@ 2 Related work
Evaluating Explanations. Current faithfulness metrics mostly use removing important tokens or retraining only on important tokens identified by the attribution methods, (Chan et al., 2022). Abnar and Zuidema (2020) use agreement with gradient and ablation methods as an evaluation of their explanation methods. Wiegreffe and Pinter (2019) acknowledges that gradient methods should not be treated as ideal or ground truth nonetheless use the gradient as a proxy of the model’s intrinsic semantics. Explanations’ trustworthiness is task- and model-specific Bastings et al. (2022), and different attribution methods give deeply inconsistent results Neely et al. (2022). hence using one explanation method as the standard or the ground truth in all scenarios does not seem to be justifiable. DeYoung et al. (2020) introduce comprehensiveness, if only the chosen important tokens are used to make the prediction (are highlighted inputs necessary), and sufficiency, if the chosen important tokens on their own are sufficient to make the prediction. Carton et al. (2020) introduce a normalized version of comprehensiveness and sufficiency by dividing these measures into null difference. The Null difference is the sufficiency of an empty input or comprehensiveness of a full input. It is unclear whether corruption techniques evaluate the informativeness of the corrupted tokens or the robustness of the model to unnatural text and the artifacts introduced during test time.
Han et al. (2020) and Jain et al. (2020) consider an attribution method to be either faithful or unfaithful, with no middle ground, and use the term ’faithful by construction.’ In contrast, other works argue that faithfulness exists on a spectrum and that we should evaluate the ’degree of faithfulness’ of an explanation method (Jacovi and Goldberg, 2020). We adopt the latter approach, searching for a sufficiently faithful explanation method for our tasks.
Atanasova et al. (2023) use counterfactuals to evaluate faithfulness of natural language explanations—When we want the model to tell us in natural language why it made a particular decision—they use counterfactual generation techniques from (Ross et al., 2021).
Another line of work tries to evaluate explanations using uncertainty estimation. Slack et al. (2021) develop a Bayesian framework for generating feature importance estimates along with their associated uncertainty in the form of credible intervals.
OOD Problem in Explainability. Hooker et al. (2019) and Vafa et al. (2021) suggest retraining or fine-tuning the model with inputs that are partly erased to achieve a model that has the same train and evaluation distribution. However, this retraining is expensive and not always possible. Kim et al. (2020) try to solve the OOD problem by making the explanation itself in-distribution.
@@
Feature Importance (Attribution). attributions are local explanations that assign a score to each input feature (token embeddings in most NLP tasks). The score represents how important that feature is for the predictor model according to that explanation method. We can categorize these methods into four types. i) Perturbation-based methods which work by perturbing the input examples such as removing or masking to measure feature importance (Li et al., 2016, 2017; Feng et al., 2018; Wu et al., 2020). ii) Gradient-Based methods determine the importance of each input feature by measuring the derivative of output with respect to each input (Mohebbi et al., 2021; Kindermans et al., 2019; Sundararajan et al., 2017; Lundstrom et al., 2022; Enguehard, 2023; Sanyal and Ren, 2021; Sikdar et al., 2021). iii) Surrogate-based methods use a simple, interpretable model to explain the original complex, black-box model (Ribeiro et al., 2016; Lundberg and Lee, 2017; Kokalj et al., 2021). iv) Decomposition-based methods try to break down the importance score into linear contributions from the input (Montavon et al., 2019; Voita et al., 2021; Chefer et al., 2021; Modarressi et al., 2022; Ferrando et al., 2022a).
@@ 3 Our method
Our faithfulness evaluation protocol consists of two models: the counterfactual generator model and the predictor model. We aim to evaluate the faithfulness of attribution methods for the predictor model. Due to the large output space of LMs, we use contrastive explanations proposed by Yin and Neubig (2022) that measure the attribution of input tokens for a contrastive model decision. Contrastive attributions try to identify the most important tokens that led the model to predict the target $y_{t}$ instead of a foil $y_{f}$. Then we use a separate editor model to change these important tokens to generate counterfactuals, i.e. generating examples that will make the original predictor model more likely to predict the foil.
The protocol to evaluate attributions consists of two phases. The first phase involves creating the editor that can generate counterfactuals. In the second phase of evaluating attributions, we use the editor and predictor together to determine what percentage of tokens the editor needs to change in order to flip the predictor model’s prediction. Figure 2 demonstrates the seconds phase.
@@
Figure 2: Our process of generating counterfactuals for evaluating attribution methods. The predictor (an LM), generates a label for the given text, and an attribution method specifics the most important tokens. We mask the top nof them and ask an editor (another LM) to change the label of the input text by filling in the masked tokens. If the attribution method is more faithful, then the needed nshould be a lower number.
Our approach for creating the editor is fine-tuning an autoregressive language model specifically for counterfactual generation. For fine-tuning, we add two tokens to the embedding space and the tokenizer, specifically ’<mask>’ and ’<counterfactual>’. For creating training examples for our counterfactual generator, inspired by Wu et al. (2021) and Donahue et al. (2020), we randomly mask between 5 and 50 percent of the tokens, then append each example’s label to it, e.g. positive or negative for SST-2 dataset, then append the <counterfactual>token, and lastly, we append the original unmasked example. The training example creation is shown in Figure 3.
In the second phase of evaluating attributions, First, we give a sentence to the predictor, then use an attribution method that was applied to the predictor to identify the most important tokens for the predictor’s decision-making process. We begin by replacing 10% of these most important tokens with ’<mask>’ and present the masked sentence and the foil label (The label with the second highest logit) to the editor to generate a counterfactual sentence (one that flips the prediction of the predictor model). If unsuccessful in flipping the prediction, we incrementally increase masking by 10% until we either flip the prediction or reach a masking threshold of 50%. This evaluation protocol is shown in Figure 2. The prompting technique used for counterfactual generation during this second phase is shown in Figure 1. The attribution technique that helps us identify the most critical tokens for creating counterfactuals and helps creating counterfactuals with the least amount of change in the original text, is the one that provides the most faithful representation of the predictor’s decision-making process.
Figure 3: Training example creation for fine-tuning the counterfactual generator, and one given sample.
@@ 4.1 Datasets
Three datasets are used for evaluating faithfulness. SST-2 (Socher et al., 2013) and IMDB (Maas et al., 2011) which are binary classification, and AG-News (Zhang et al., 2015) a four class classification dataset.
Faithfulness evaluation datasets should not have gold attribution labels because we do not want human intuition in faithfulness evaluation. We want to know how the model makes a prediction (Jacovi and Goldberg, 2020).
@@ 4.2.1 Editor Models
For the editor model our method is very similar to Wu et al. (2021) which uses GPT-2 which is a decoder-only causal model for generating counterfactuals. We use three more modern decoder-only models. GPT-J-6B Wang and Komatsuzaki (2021) which we call gptj , and two sizes of Pythia. Pythia-1.4B (pythia1), and Pythia-2.8B (pythia2) Biderman et al. (2023). We fine-tune these models in accordance with Section 3. pythia1 model is fully fine-tuned but the other two are tuned using Low Rank Adaptation (LoRA) (Hu et al., 2022). We train these models for 8 epochs using dynamic masking Liu et al. (2019), which means masking each example differently in each epoch.
@@ 4.2.2 Predictor Models
We use Gemma-2b Team et al. (2024) as the predictor. We fine-tune the raw language model for the three tasks (gemma-ft). We use LoRA for fine-tuning. We also use an off-the-shelf instruct-tuned version (gemma-it) in zero-shot scenario. We compare these two versions (fine-tuned and non-fine-tuned) extensively for attribution evaluation.
@@ 4.3 Attribution Methods
Here we detail the six attribution methods that we use which are all widely used methods. We use all attribution methods in a contrastive way (Yin and Neubig, 2022). Contrastive attributions measure which features from the input make the foil token $y_{f}$ more likely and the target token $y_{t}$ less likely. We denote contrastive, target, and foil attributions by $S^{C}$, $S^{t}$, and $S^{f}$ respectively:
@@
We use implementation of these attribution methods provided by Yin and Neubig (2022) (for Gradient $\times$ input, gradient norm and erasure) and by Captum Miglani et al. (2023) (for KernelSHAP and Integrated Gradient).
@@ 4.3.1 Gradient Norm
We can calculate attributions based on the norm of the gradient of the model prediction, with respect to the input (Simonyan et al., 2013; Li et al., 2016). Gradient with respect to feature $x_{i}$:
@@
We will use both norm one and norm two:

### v3

@@ Abstract
Despite the widespread adoption of autoregressive language models, explainability evaluation research has predominantly focused on span infilling and masked language models. Evaluating the faithfulness of an explanation method—how accurately it explains the inner workings and decision-making of the model—is challenging because it is difficult to separate the model from its explanation. Most faithfulness evaluation techniques corrupt or remove input tokens deemed important by a particular attribution (feature importance) method and observe the resulting change in the model’s output. However, for autoregressive language models, this approach creates out-of-distribution inputs due to their next-token prediction training objective. In this study, we propose a technique that leverages counterfactual generation to evaluate the faithfulness of attribution methods for autoregressive language models. Our technique generates fluent, in-distribution counterfactuals, making the evaluation protocol more reliable.
@@ 1 Introduction
Most modern NLP systems rely on autoregressive, transformer-based language models (Brown et al., 2020; Touvron et al., 2023; Groeneveld et al., 2024). These models are inherently opaque, creating a strong need to understand their decision-making processes. As a result, explanation methods have become increasingly important in the field.
A widely-used approach for model explainability is attribution, also known as feature importance (FI) (Zhao et al., 2023). Attribution methods aim to identify which input features contribute most to a model’s predictions, assigning a scalar value to each feature that reflects its relevance in the decision-making process. In typical NLP tasks, input features are often subwords or their combinations.
@@
A key challenge in evaluating the faithfulness of attribution methods is that many existing techniques are designed for denoising or masked language models (MLMs) (Kobayashi et al., 2020, 2021; Ferrando et al., 2022b; Modarressi et al., 2022, 2023; Mohebbi et al., 2023). Recent work on autoregressive models has primarily focused on the plausibility of attributions (Yin and Neubig, 2022; Ferrando et al., 2023). While plausible (or persuasive) explanations might be the objective of the explainer, the core objective for the user is to truly understand the model’s decision-making process, rather than simply being convinced that the model’s decisions are correct (Jacovi and Goldberg, 2021).
Nearly all previous methods for faithfulness evaluation modify the input in some way, such as masking or removing important tokens based on the attribution results, and then measuring the impact on the model’s predictions. These methods tend to work well for MLMs, which are specifically trained for tasks like span or mask infilling. However, in the case of autoregressive models like GPT-2, which predict the next token, such modifications produce out-of-distribution (OOD) inputs. This raises a crucial question: are these evaluation methods truly assessing the informativeness of the selected tokens, or merely testing the model’s robustness to unnatural text and the artifacts introduced by testing modifications (Hooker et al., 2019)? Moreover, the OOD nature of these inputs results in explanations that become socially misaligned (Hase et al., 2021). In other words, the expectations of users—who seek to understand which features are most relevant to the model’s decision—no longer align with the actual output of the attribution method. Instead, feature importance becomes influenced by the model’s priors rather than the learned features that truly drive predictions.
In this work, drawing inspiration from counterfactual generation—where the input is altered to flip the model’s output—we propose a new technique to evaluate the faithfulness of attribution methods in autoregressive language models. Specifically, we use counterfactual generators to modify the input by focusing on tokens highlighted by attribution methods, while ensuring that the altered input remains natural, fluent, and within the model’s original distribution. This ensures that any observed change in the model’s predictions is due to the modification of the important tokens, rather than an effect of OOD inputs.
@@
We argue that if an attribution method enables a counterfactual generator to modify fewer tokens to change the model’s prediction, then it demonstrates a stronger understanding of the model’s inner workings, indicating higher faithfulness. To validate our approach, we apply this faithfulness evaluation technique to several attribution methods—including gradient norm, gradient × input, erasure, KernelSHAP, and integrated gradients—within the context of next-word prediction for two language models: the fine-tuned Gemma-2b and the off-the-shelf Gemma-2b-instruct (Team et al., 2024).
@@
Our contributions are as follows: (i) We introduce a novel faithfulness evaluation protocol that preserves the model’s input distribution, designed for attribution methods in autoregressive language models. (ii) We apply this protocol to evaluate and rank widely-used attribution methods, showcasing differences in sensitivity between fine-tuned and off-the-shelf models when handling OOD data and proposing a solution. [^1] The code is available at https://github.com/Sepehr-Kamahi/faith
@@ 2 Related work
Evaluating Explanations. Most current metrics for evaluating faithfulness involve either removing important tokens or retraining the model using only those identified as important by attribution methods (Chan et al., 2022). For instance, Abnar and Zuidema (2020) assess explanations by comparing them with gradient and ablation techniques. Although Wiegreffe and Pinter (2019) caution that gradients should not be considered ideal or the “ground truth,” they still utilize gradients as a proxy for the model’s intrinsic semantics. Importantly, the trustworthiness of explanations is both task- and model-dependent (Bastings et al., 2022), and different attribution methods frequently produce inconsistent results (Neely et al., 2022). As a result, it is not justifiable to treat any single explanation method as a universal standard across all contexts.
In their work, DeYoung et al. (2020) introduce two key concepts: comprehensiveness (whether the important tokens identified are the only ones necessary for making a prediction) and sufficiency (whether these important tokens alone are enough to make the prediction). Carton et al. (2020) build on this by proposing normalized versions of these concepts, comparing comprehensiveness and sufficiency to the null difference—the performance of an empty input (for sufficiency) or a full input (for comprehensiveness). However, it remains unclear whether these corruption techniques evaluate the informativeness of the corrupted tokens or merely the robustness of the model to unnatural inputs and artifacts introduced during evaluation.
Further, Han et al. (2020) and Jain et al. (2020) frame attribution methods as either faithful or unfaithful, with no consideration for degrees of faithfulness. They describe attribution methods that are "faithful by construction." In contrast, other researchers propose that faithfulness exists on a spectrum and suggest evaluating the “degree of faithfulness” of explanation methods (Jacovi and Goldberg, 2020). Our approach aligns with this view, as we aim to find explanation methods that are sufficiently faithful for autoregressive models.
Atanasova et al. (2023) evaluate the faithfulness of natural language explanations using counterfactuals, applying techniques from Ross et al. (2021) to assess how well explanations align with the model’s decision-making. This line of work offers valuable insights into the use of counterfactuals, which we build upon for evaluating attribution methods in language models. Another relevant direction is the evaluation of explanations using uncertainty estimation. For example, Slack et al. (2021) develop a Bayesian framework that generates feature importance estimates along with their associated uncertainty, expressed through credible intervals, highlighting the importance of uncertainty in faithfulness evaluations
The OOD Problem in Explainability.
The issue of OOD inputs in explainability has been raised by several works. Hooker et al. (2019) and Vafa et al. (2021) suggest retraining or fine-tuning the model using partially erased inputs to align training and evaluation distributions. However, this process can be computationally expensive and is not always practical. An alternative approach by Kim et al. (2020) aims to ensure that the explanation remains in-distribution to mitigate OOD problems. Our work addresses this concern by preserving the input distribution during faithfulness evaluation, particularly for autoregressive models.
@@
Feature Importance (Attribution). Attributions, or feature importance scores, are local explanations that assign a score to each input feature—typically token embeddings in NLP tasks—indicating how crucial that feature is to the model’s prediction. Attribution methods can be categorized into four types: i) Perturbation-based methods, which alter or mask input features to assess their importance by observing changes in the model’s output (Li et al., 2016, 2017; Feng et al., 2018; Wu et al., 2020). ii) Gradient-based methods, which calculate the derivative of the model’s output with respect to each input to measure the influence of each feature (Mohebbi et al., 2021; Kindermans et al., 2019; Sundararajan et al., 2017; Lundstrom et al., 2022; Enguehard, 2023; Sanyal and Ren, 2021; Sikdar et al., 2021). iii) Surrogate-based methods, which explain a complex black-box model using a simpler, interpretable model (Ribeiro et al., 2016; Lundberg and Lee, 2017; Kokalj et al., 2021). iv) Decomposition-based methods, which break down the overall importance score into linear contributions from the input features (Montavon et al., 2019; Voita et al., 2021; Chefer et al., 2021; Modarressi et al., 2022; Ferrando et al., 2022a).
@@ 3 Our method
Our faithfulness evaluation protocol consists of two models: a counterfactual generator and a predictor model. Our goal is to assess the faithfulness of attribution methods used for the predictor model. Given the large output space of autoregressive language models (LMs)—often encompassing thousands of vocabulary items—analyzing the entire output space provides limited insight. To address this, we adopt contrastive explanations from Yin and Neubig (2022), which evaluate the attribution of input tokens for a contrastive decision. Specifically, contrastive attributions identify the most influential tokens that led the model to predict the target label $y_{t}$ over an alternative foil $y_{f}$. These important tokens are then modified by a separate editor model to generate counterfactuals, i.e., examples that shift the prediction of the original model towards the foil.
Our evaluation protocol for attributions involves two key phases. In the first phase, we create an editor capable of generating counterfactuals. The second phase involves using both the editor and predictor to determine the minimum percentage of tokens the editor must change to flip the predictor’s prediction. Figure 2 illustrates the second phase of this process.
To construct the editor, we fine-tune an autoregressive language model for the task of counterfactual generation. During fine-tuning, we expand the model’s embedding space and tokenizer by adding two special tokens: <mask>’ and <counterfactual>’. Inspired by Wu et al. (2021) and Donahue et al. (2020), we generate training examples by randomly masking between 5% and 50% of the tokens in each example. To each example, we append its label (e.g., positive or negative for the SST-2 dataset), the ‘<counterfactual>’ token, and the original unmasked example. Figure 3 demonstrates this process of constructing the training examples.
In the second phase of attribution evaluation, we input a sentence into the predictor model and apply an attribution method to identify the most important tokens influencing its prediction. Starting with the top 10% most important tokens, we replace these tokens with ‘<mask>’. The masked sentence, along with the foil label (the label associated with the second-highest logit), is then passed to the editor, which generates a counterfactual sentence intended to flip the predictor model’s output. If the prediction does not flip, we incrementally increase the masking by 10% until either the prediction is flipped or we reach a masking threshold of 50%. This process is outlined in Figure 2, and the prompting technique for counterfactual generation is detailed in Figure 1. The attribution method that enables the editor to flip the prediction with the fewest changes is deemed to provide the most faithful representation of the predictor model’s decision-making process.
@@
Figure 2: Our process of generating counterfactuals for evaluating attribution methods. The predictor (an LM) generates a label for the given text, and an attribution method specifies the most important tokens. We mask the top nof them and ask an editor (another LM) to change the label of the input text by filling in the masked tokens. If the attribution method is more faithful, then the required nshould be lower.
Figure 3: Creation of training examples for fine-tuning the counterfactual generator, and one given sample.
@@ 4.1 Datasets
We use three datasets for evaluating faithfulness: SST-2 (Socher et al., 2013) and IMDB (Maas et al., 2011), which are both binary classification datasets, and AG-News (Zhang et al., 2015), a four-class classification dataset.
Faithfulness evaluation datasets should not have gold attribution labels because we do not want human intuition to influence the evaluation. Instead, we aim to understand how the model makes predictions (Jacovi and Goldberg, 2020).
@@ 4.2.1 Editor Models
For the editor model, our method is similar to Wu et al. (2021), which uses GPT-2, a decoder-only causal model, for generating counterfactuals. We extend this by using three more modern decoder-only models: GPT-J-6B (Wang and Komatsuzaki, 2021), which we refer to as "gptj," and two sizes of Pythia: Pythia-1.4B (pythia1) and Pythia-2.8B (pythia2) Biderman et al. (2023). We fine-tune these models following the process described in Section 3. The pythia1 model is fully fine-tuned, while the other two (gptj and pythia2) are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2022). All models are trained for 8 epochs using dynamic masking (Liu et al., 2019), meaning each example is masked differently in each epoch.
@@ 4.2.2 Predictor Models
We use Gemma-2b (Team et al., 2024) as the predictor model. We fine-tune the raw language model for the three datasets (referred to as gemma-ft) using Low-Rank Adaptation (LoRA). Additionally, we employ an off-the-shelf instruct-tuned version (gemma-it) for zero-shot evaluation. We then conduct a detailed comparison between these two versions—fine-tuned (gemma-ft) and non-fine-tuned (gemma-it)—to assess their differences in attribution evaluation.
@@ 4.3 Attribution Methods
Here we detail the six widely used attribution methods employed in our study. We use all attribution methods in a contrastive way (Yin and Neubig, 2022). Contrastive attributions measure which features from the input make the foil token $y_{f}$ more likely and the target token $y_{t}$ less likely. We denote contrastive, target, and foil attributions by $S^{C}$, $S^{t}$, and $S^{f}$ respectively:
@@
We use the implementation of these attribution methods provided by Yin and Neubig (2022) (for Gradient $\times$ input, gradient norm and erasure) and by Captum (Miglani et al., 2023) (for KernelSHAP and Integrated Gradient).
@@ 4.3.1 Gradient Norm
We can calculate attributions based on the norm of the gradient of the model’s prediction with respect to the input $x$ (Simonyan et al., 2013; Li et al., 2016).The gradient with respect to feature $x_{i}$ is given by:
@@
We will use both norm one (gradnorm1) and norm two (gradnorm2):